ROTATIONALLY RESOLVED VIBRATIONAL OVERTONE SPECTRA OF JET-COOLED METHANOL

OLEG V. BOYARKIN, LUCIA LUBICH, \underline{THOMAS R. RIZZO}, Laboratoire de chimie physique moléculaire, Ecole Polytechnique Fédérale de Lausanne, CH- Ecublens, 1015 Lausanne, Switzerland; and DAVID S. PERRY, Department of Chemistry, University of Akron, Akron, Ohio, 44325.

Infrared-optical double resonance excitation permits the measurement of fully rotationally resolved vibrational overtone spectra of jet-cooled methanol up to the 6ν_{OH} level. These weak transitions to high vibrational levels are detected using the technique of IR assisted photofragment spectroscopy (IRLAPS). The elimination of spectral congestion reveals splittings of the zeroth-order rovibrational levels that reflect vibrational coupling to dark background levels. Analysis of the overtone spectra as a function of the number of OH stretch quanta indicates the presence of a strong coupling between zeroth-order states with nν_{OH} and those with (n-1)ν_{OH} + ν_{CHasym.st.}. This coupling is most clearly manifest at the 5ν_{OH} level where the OH stretch frequency is anharmonically shifted close to that of the CH asymmetric stretch frequency. Finer spectral splittings reflect the coupling of the mixed OH-CH stretch states with the rest of the vibrational bath. The most recent results using this technique for CH_{3}OH and its partially deuterated analogs will be discussed.